OpenStudy (anonymous):

(2y^2-7y-15)/(3y^2-8y-3) times (9y^2-1)/(4y^2-9) divided by (y^2+3y-10)/(2y^2-9y+9) help??

OpenStudy (amistre64):

\[\frac{(2y^2-7y-15)}{(3y^2-8y-3)}.\frac{(9y^2-1)}{(4y^2-9)} \div \frac{(y^2+3y-10)}{(2y^2-9y+9)}\] flip the last one over; factor it all; cancel the like terms top and bottom and simplify :)

OpenStudy (amistre64):

i see a few difference of squares that factor nicely

OpenStudy (anonymous):

Okay! Lets see what I get... hahaha

OpenStudy (amistre64):

(2y+3)(y-5)(3y+1)(3y-1)(2y-3)(y-3) ------------------------------- (y-3)(3y+1)(2y-3)(2y+3)(y-5)(y+2) (3y-1) ----- is what I get (y+2)

OpenStudy (anonymous):

After factoring and cancelling I got.... (3y-1)(y-5)/(y+5)(y-2) I think I might have messed up though in the middle :/

OpenStudy (amistre64):

its alot to keep track of :) but if you take it step by step; and verify your factors; your answer should match mine :)

OpenStudy (anonymous):

okay! Thanks :)

OpenStudy (amistre64):

lol ... or mine will match yours :)

OpenStudy (anonymous):

hahaha :)

OpenStudy (amistre64):

(y^2+3y-10) (y+5)(y-2) is right; yours wins lol

OpenStudy (anonymous):

really?? :D

OpenStudy (amistre64):

(2y^2-7y-15) (y-5)(2y+3); yep; yours is good ;)

OpenStudy (anonymous):

yay!! thank you!! :)